# Guarantees in Security: A Philosophical Perspective

### v4

@@ I Introduction
In the Pwn2Own hacking contest this year, a single person (Manfred Paul) demonstrated successful exploits for all major browsers: Google Chrome, Mozilla Firefox, Apple Safari, and Microsoft Edge [1]. Just Chrome is used by 3.5 billion people. [^1] https://backlinko.com/chrome-users Today, thanks to Paul, these major security flaws are fixed. Last year, evidently, most people on Earth accessed the internet with the door open to potential attackers—despite decades of research in defensive security, despite an abundance of mitigations, despite a dedicated red team (Google P0) that is considered the best in the world. From a software engineering perspective, research in cybersecurity might seem technology-specific, reactive, and ostensibly ineffective. Like a snake eating its own tail, offense and defense forever on repeat.
But why? Software is entirely virtual and can be described completely [2]: A program’s source code is meant to formally express the programmer’s intention using the syntactic and semantic rules of the programming language. As the behavior of a software system arises from well-defined instructions, we must be able to formally reason about all its properties. Surely there exists an approach that will forever guarantee the security of our systems. Only, we haven’t found it, yet?
@@
While an important question for our ICSE community, our battle-tested research methods are not well equipped to answer it. Since we do not propose a new approach or intervention, we can hardly run measurements or experiments to gain insight. Since we look for guidance on the science of security itself, we cannot expect answers from descriptive methods, either.
In this paper, we take a philosophical perspective on guarantees in cybersecurity and develop a first theory of cybersecurity that explains why we must, in general, assume that exploitable security flaws exist in any sufficiently large software system. We substantiate each argument by demonstrating how the corresponding challenge is routinely exploited to attack a system despite credible assurances about the absence of security flaws. We believe, now is the time to reflect on our scientific progress more fundamentally and to shed light on the limitations of our approaches to reason about cybersecurity. We do not claim, our theory is final. Instead, we expect our theory to be discussed, improved, and refined in a protracted scientific debate.
What does this mean for cybersecurity as a research field? The quest for truth in the absence of a mechanism, guaranteed to arrive at the truth, has existed since the time of enlightment. In his philosophy of science, Popper suggests that a scientific theory must evolve in a counterexample-guided manner via falsification [3]. In his philosophy of maths, Lakatos suggests that even mathematical theorems evolve in a counterexample-guided manner via proofs and refutations [4]. Similarly, we suggest in our philosophy of security that claims of security must evolve in a counterexample-guided manner, too.
Consider fishing as a metaphor for bug finding where a fishernet represents our cybersecurity tools and processes while the fishes represent the security flaws in our software systems. Our theory explains why, for every given net, there will always be a fish that slips through the net. Clearly, this does not undermine the utility of the fishernet. Our philosophy guides the perspective from the need to develop the ultimate fishernet to an incremental, counterexample-guided evolution of our nets to approach maximum effectiveness empirically.
@@ II-B Generating Reliable Statements about Software Security
1)  What is a secure software software system?
@@ III-A Challenges to Distinguish Secure from Insecure Software
Example 1. For instance, Nakamoto’s consensus protocol has been formally shown to “guarantee safety and liveness” [22]. However, there is nothing preventing malicious participants from violating their part of that protocol and casting their vote more than once. To account for this, we need a secure consensus protocol to have another property called accountable safety [23]. Moreover, there is nothing preventing an attacker from creating network partitions such that the block confirmation procedure cannot be relied upon. To guarantee safety under network partitions, we need a secure consensus protocol to have another property called finality [24]. Many other properties need to be satisfied for a consensus protocol to be able to withstand attacks, some of which we are known while others remain unknown until their absence is exploited.
@@ III-C Sociotechnical Challenges
Example 20. Project maintainers or owners rely on a code review process to check the absence of bugs in new contributions. While we can safely assume that the vast majority of contributions are benign, without additional tooling there are opportunities for an attacker to exploit the developer’s trust in code contributions and inject vulnerabilities that could later be exploited [69]. [^17] XZ open source attack (Jia Tan): https://research.swtch.com/xz-timeline
@@
  • Even if we trust the security policies of the linux kernel, there may be bugs in the eBPF verifier181818eBPF (https://eBPF.io) enables users to instrument a running system by loading small programs into the operating system kernel. It is used, e.g., to implement security policies. preventing us from detecting policy violations [70, 71, 72].
@@
Problem of Infinite Regress [76]. In epistemic philosophy, we call this the problem of Infinite Regress: A belief is justified because it is based on another belief that is justified. We must anchor our trust somewhere. While these trust anchors are often well-specified in security, they are not exempt from scrutiny and subject to the same challenges we have listed in support of our theory so far. An attacker may be able to attack the software system by exploiting bugs or security flaws in our trust anchors.
@@
Example 22. While the developer takes a constructive perspective, the attacker takes an adversarial perspective [21]. On the one hand, the developer reads code to implement the intended features and to fix bugs. On the other hand, the attacker reads code (or binary) to exploit the unintented features (i.e., the weird machine) [61, 62, 36]. These conflicting perspectives must be reconciled. The developer implements with intention but might realize something that deviates from this intention [77]. This difference between intention and realization can be exploited “to program the weird machine” [62].
@@
Mitigations as cost. Any mechanism or process to improve the security of a software system also increases the cost which subtracts from the incentive. Due to this incentive structure, an attacker will always target low hanging fruits first. Once a mitigation is introduced and adopted to prevent a large class of security flaws currently exploited, attackers will just move on to the next low hanging fruit [83].
Security-by-obscurity not viable. According to Kerckhoff’s principle, a system should be secure even if everything about the system is public knowledge. Obfuscation and other deterrents increase the imbalance between attacker and defender in favor of the attacker. While there is obviously no incentive for ethical hackers to discover and responsibly disclose existing security flaws, a determined attacker can be expected to have more skills, resources, and incentive to exploit the (unhandled) security flaws despite the deterrent. The security of a system should not depend on the secrecy of its components. The reputational and financial damage arising from unknown exploits of security flaws is significantly worse than inviting responsible disclosure via an open security approach.
Example 25. Dellago et al. [45] studied the market for 0-day exploits from 2016 and 2021.2020200-day exploits use vulnerabilities that are unknown to the developers. For the four studied operating systems (Windows, MacOS, iOS, and Android), the number of 0-days observed in the wild decreased from 2016 to 2018 but increased again until the end of the study period. For the two investigated exploit brokers, Zerodium and Crowdfence, exploit prices increased to a maximum in 2019 and have since remained stable at USD 2.5-3 million for a full exploit chain with persistence on Android or iOS. Recently, a third broker, Operation Zero, has announced to offer USD 20 million for the same, which Google’s Senior Director of the Threat Analysis Group, Shane Huntley, took as a good sign that the developed mitigations are becoming more effective and exploits more difficult. [^21] https://twitter.com/ShaneHuntley/status/1706944206521160070 Schechter [84] confirms that higher prices are a signal of more secure products.
@@ Acknowledgments
This paper was written out of a deep curiousity to understand the reliability of the statements that our approaches can generate about the properties of a software system, and how we should proceed if we accepted that these statements are fundamentally unreliable from a formal point of view. On the one hand, researchers perpetually promise guarantees and effectiveness. On the other hand, practitioners are troubled by the obvious lack of guarantees. This paper is an attempt to shed some first light and to reconcile the promise with reality.
Without revealing our identity, we extend our special thanks to Toby Murray, Joshua J. Drake, Daniel Woods, Nathanial (d0nut) Lattimer, and Thomas (Halvar Flake) Dullien for the insightful feedback and inspiring discussions.

### v5

# Fundamental Challenges in Cybersecurity and a Philosophy of Vulnerability-Guided Hardening

@@ I Introduction
In the Pwn2Own hacking contest this year, a single person (Manfred Paul) demonstrated successful exploits for all major browsers: Google Chrome, Mozilla Firefox, Apple Safari, and Microsoft Edge [1]. Just Chrome is used by 3.5 billion people. [^1] https://backlinko.com/chrome-users Today, thanks to Paul, these major security flaws are fixed. Last year, evidently, most people accessed the internet with the door open to potential attackers—despite decades of research in defensive security, despite an abundance of mitigations, despite a dedicated red team (Google P0) that is considered the best in the world. From the outside, research in cybersecurity might seem technology-specific, reactive, and ostensibly ineffective.
But why? Software is entirely virtual and can be described completely [2]: A program’s source code is meant to formally express the programmer’s intention using the syntactic and semantic rules of the programming language. As the behavior of a software system arises from well-defined instructions, we must be able to formally reason about all its properties. Surely there exists an approach that will forever guarantee the security of our systems. Only, we have not found it, yet?
@@
This question about the effectiveness of program analysis is at the very heart of Software Engineering. Yet, our battle-tested research methods are not well equipped to answer it. Since we do not propose a new approach or intervention, we can hardly run measurements or experiments to gain insight. Since we look for guidance on the science of security itself, we cannot expect answers from descriptive methods, either.
In this paper, we take an objective, philosophical perspective on the effectiveness of cybersecurity tooling and develop a theory of cybersecurity that explains why we must, in general, assume that exploitable security flaws exist in any sufficiently large software system. We substantiate each argument by demonstrating how the corresponding challenge is routinely exploited to attack a system despite credible assurances about the absence of security flaws. We believe, now is the time to reflect on our scientific progress more fundamentally and to shed light on the limitations of our approaches to reason about the properties of a system. We do not claim, our theory is final or complete. Instead, we expect it to be discussed, improved, and refined in a protracted scientific debate.
How do we make meaningful progress? The quest for truth in the absence of a mechanism guaranteed to arrive at the truth has existed since the time of enlightment. In his philosophy of science, Popper suggests, scientific theories evolve in a counterexample-guided manner (conjectures and refutations[3]. In his philosophy of maths, Lakatos suggests, even theorems evolve in a counterexample-guided way (proofs and refutations[4]. Similarly, in our philosophy of security, we suggest, security claims evolve via counterexamples, too.
Consider fishing as a metaphor for bug finding. A fishernet represents our tools and processes while the fishes represent the security flaws in our software systems. Our theory explains why, for every net, there will always be a fish that slips through. Clearly, this does not undermine the utility of the net. Our philosophy turns the objective from developing the ultimate fishernet to an incremental, counterexample-guided evolution of our nets maximize effectiveness empirically.
@@ II-B Generating Reliable Statements about Software Security
1)  What is a secure software system?
@@ III-A Challenges to Distinguish Secure from Insecure Software
Example 1. For instance, Nakamoto’s consensus protocol has been formally shown to “guarantee safety and liveness” [22]. However, there is nothing preventing malicious participants from violating their part of that protocol and casting their vote more than once. To account for this, we need a secure consensus protocol to have another property called accountable safety [23]. Moreover, there is nothing preventing an attacker from creating network partitions such that the block confirmation procedure cannot be relied upon. To guarantee safety under network partitions, we need a secure consensus protocol to have another property called finality [24]. Many other properties need to be satisfied for a consensus protocol to be able to withstand attacks, some of which are known while others remain unknown until their absence is exploited.
@@ III-C Sociotechnical Challenges
Example 19. Cryptography ensures the confidentiality of some data (e.g., messages) only if we trust the confidentiality of other data (e.g., secret keys). However, secret keys can be compromised. For instance, Intel SGX-based trusted execution environments have recently been compromised because the root provisioning key (SGX Fuse Key0) and the root sealing key (SGX Fuse Key1) can be leaked. [^17] https://x.com/PratyushRT/status/1828183761055330373
@@
Example 21. Project maintainers rely on a code review process to check the absence of bugs in new contributions. While we can safely assume that the vast majority of contributions are benign, without additional tooling there are opportunities for an attacker to exploit the developer’s trust in code contributions and inject vulnerabilities that could later be exploited [69]. [^18] XZ open source attack (Jia Tan): https://research.swtch.com/xz-timeline
@@
  • Even if we trust the security policies of the linux kernel, there may be bugs in the eBPF verifier191919eBPF (https://eBPF.io) enables users to instrument a running system by loading small programs into the operating system kernel. It is used, e.g., to implement security policies. preventing us from detecting policy violations [70, 71, 72].
@@
Problem of Infinite Regress [76]. In epistemic philosophy, this is the problem of Infinite Regress: A belief is justified because it is based on another belief that is justified. We must anchor our trust somewhere. While these trust anchors are often well-specified in security, they are not exempt from scrutiny and subject to the same challenges we have listed in support of our theory. An attacker may be able to attack the system by exploiting bugs or security flaws in our trust anchors.
@@
Example 23. While the developer takes a constructive perspective, the attacker takes an adversarial perspective [21]. On the one hand, the developer reads code to implement intended features and to fix bugs. On the other hand, the attacker reads code (or binary) to exploit unintented features (i.e., the weird machine) [61, 62, 36]. These conflicting perspectives must be reconciled. The developer implements with intention but might realize something that deviates from this intention [77]. This difference between intention and realization can be exploited “to program the weird machine” [62].
Example 24. Better performance often wins against better security. For instance, the browser security team of Microsoft Edge recently found that disabling the Just-In-Time (JIT) optimization could mitigate half of the Chrome exploits observed in-the-wild. [^20] https://microsoftedge.github.io/edgevr/posts/Super-Duper-Secure-Mode/ In fact, performance optimization is a root cause of many critical types of security flaws. For instance, memory corruption (such as buffer overflow or use-after-free) is often caused by undefined behavior which is deliberately left undefined in the C language specification for optimization reasons. Side channel attacks are often enabled by microarchitectural or compiler optimizations. For some critical types of security flaws, we even have mitigations. However, for performance reasons they are often left disabled in production.
@@
Example 26. Some maintainers remain unconvinced that rewriting components of their project in a memory-safe langauge will help improve the security of the project:
@@
Mitigations as cost. Any mechanism or process to improve the security of a system also increases the cost which subtracts from the incentive. Due to this incentive structure, an attacker will always target low hanging fruits first. Once a mitigation is adopted to prevent a vulnerability class currently popular, attackers will just move on to the next low hanging fruit [83].
Security-by-obscurity not viable. According to Kerckhoff’s principle, a system should be secure even if everything about the system is public knowledge. Obfuscation and other deterrents increase the imbalance between attacker and defender in favor of the attacker. While there is obviously no incentive for ethical hackers to discover and responsibly disclose existing security flaws, a determined attacker can be expected to have more skills, resources, and incentive to exploit the (unhandled) security flaws despite the deterrent. The security of a system should not depend on the secrecy of its components. The reputational and financial damage arising from unknown exploits is significantly worse than inviting responsible disclosure via an open security approach.
Example 28. Dellago et al. [45] studied the market for 0-day exploits from 2016 and 2021.2222220-day exploits use vulnerabilities that are unknown to the developers. For the four studied operating systems (Windows, MacOS, iOS, and Android), the number of 0-days observed in the wild decreased from 2016 to 2018 but increased again until the end of the study period. For the two investigated exploit brokers, Zerodium and Crowdfence, exploit prices increased to a maximum in 2019 and have since remained stable at USD 2.5-3 million for a full exploit chain with persistence on Android or iOS. Recently, a third broker, Operation Zero, has announced to offer USD 20 million for the same, which Google’s Senior Director of the Threat Analysis Group, Shane Huntley, took as a good sign that the developed mitigations are becoming more effective and exploits more difficult. [^23] https://twitter.com/ShaneHuntley/status/1706944206521160070 Schechter [84] confirms that higher prices are a signal of more secure products.
@@ Acknowledgments
This paper was written out of a deep curiousity to understand the reliability of the statements that our approaches can generate about the properties of a software system, and how we should proceed if we accepted that these statements are fundamentally unreliable from a formal point of view. On the one hand, researchers perpetually promise guarantees and effectiveness. On the other hand, practitioners are troubled by the obvious lack of guarantees. This paper is an attempt to reconcile the promise with reality: Even a formal security proof must end with the invitation to find new vulnerabilities.
We extend our special thanks to Toby Murray, Joshua J. Drake, Daniel Woods, Nathanial (d0nut) Lattimer, and Thomas (Halvar Flake) Dullien for the insightful feedback and inspiring discussions.